# arXiv:2603.00822v1[cs.SE] 28 Feb 2026

### 28 Feb 2026 v1

@@ 5. Discussion
The shift from human-supervised to autonomous agents fundamentally changes the role of project documentation. Context Drift, where agents produce changes that violate their instructions, is not merely a technical bug but a symptom of a deeper mismatch between how we document intent and how agents consume it. We discuss the theoretical and practical implications of treating Agent Instructions as executable specifications to bridge this Reality Gap.
@@ 5.2. The Agentic Feedback Loop
When an agent hits a ContextCov enforcement block, the error is fed back into its context window. This empowers the agent to reason about the failure. If the agent determines the rule was poorly scoped, it can autonomously propose a fix to the Agent Instructions. ContextCov transforms the agent from a passive consumer of documentation into an active maintainer of project specifications. Our design adopts a strict, fail-closed interpretation of documented constraints (Section 3.5). We argue that ambiguity in Agent Instructions is itself a defect. If a rule has valid exceptions, those exceptions should be specified explicitly. By blocking ambiguous cases, ContextCov forces Agent Instructions to become more precise over time, progressively narrowing the Reality Gap.
@@ 5.3. Executable Interpretability
It is important to distinguish ContextCov from security sandboxes. Security mechanisms like SELinux (SELinux Project, ), AppArmor (AppArmor, ), and seccomp (Linux Kernel, ) enforce mandatory access controls at the kernel level, while container runtimes like gVisor (Google, ) and Firecracker (Amazon Web Services, ) provide process isolation. Runtime verification frameworks (Havelund and Rosu, 2001; Chen and Roşu, 2007; Runtime Verification Inc., ) monitor program execution against formal specifications. ContextCov is not designed to prevent a malicious actor from compromising a system. Rather, it is designed to prevent a well-intentioned but hallucinating agent (Lin et al., 2024) from degrading the repository. By restricting the execution environment through mechanisms like the Process Interceptor, ContextCov bounds the agent’s action space, ensuring that its exploratory problem-solving stays within approved architectural and operational limits.
@@ 6. Related Work
ContextCov draws on and extends several lines of prior research.
Documentation-to-Code Consistency.
@@
Prior work has studied the “stale docs” problem (Aghajani et al., 2019), API documentation issues (Uddin and Robillard, 2015; Robillard and DeLine, 2011; Treude and Robillard, 2016), and README files (Prana et al., 2018; Gaughan et al., 2025). Techniques detect comment-code inconsistencies (Tan et al., 2007, 2012; Ratol and Robillard, 2017) and update comments after code changes (Panthaplackel et al., 2020; Liu et al., 2020). Requirements traceability research (Gotel and Finkelstein, 1994; Antoniol et al., 2002) maintains links between specifications and implementations. However, prior work focuses on descriptive documentation (docstrings, comments) for human developers. ContextCov targets prescriptive documentation (Agent Instructions) where stale docs become execution hazards since agents treat instructions literally.
Natural Language to Specifications and Linters.
@@
Research has translated natural language to formal specifications (Ghosh et al., 2016; Brunello et al., 2019; Pandita et al., 2012; Zhong et al., 2009). Specification mining extracts properties from traces (Ammons et al., 2002; Gabel and Su, 2008; Kremenek et al., 2006; Ernst et al., 2007), while learning-based approaches infer linting rules (Bader et al., 2019). Tools like ESLint (21), Pylint (48), Semgrep (27), Infer (Calcagno et al., 2015), and Error Prone (Aftandilian et al., 2012) enforce standards but require manual configuration or focus on general defects. ContextCov generates checks from Agent Instructions using domain-routed synthesis, leveraging code search techniques (Di Grazia and Pradel, 2023) to produce unified Python scripts operating on Tree-sitter ASTs, NetworkX graphs, and shell shims.
LLM Agents and Alignment.
@@
Research on coding agents (Yang et al., 2024; Wang et al., 2025) has demonstrated autonomous repository navigation and issue resolution, with benchmarks like SWE-bench (Jimenez et al., 2024) and BigCodeBench (Zhuo et al., 2024) driving rapid progress. Agent alignment frameworks like Constitutional AI (Bai et al., 2022) and Guardrails (Guardrails AI, ) focus on general-purpose safety constraints, while research on jailbreaking (Wei et al., 2024) reveals vulnerabilities in instruction-following. ContextCov provides domain-specific alignment for software engineering contexts, ensuring agents adhere to project-level conventions rather than just general safety guidelines. While Constitutional AI constrains what an LLM may generate, ContextCov constrains what an agent may execute in a specific repository context.
Architectural Conformance.
@@
Tools like Sonargraph (33) detect dependency violations, while reflexion models (Murphy et al., 2001) compare intended and actual architectures. Knodel and Popescu (Knodel and Daniel, 2007) survey architecture compliance checking approaches, and ArchJava (Chambers and Notkin, 2002) embeds architectural constraints directly in the programming language. These tools require architects to manually encode rules in tool-specific formats. ContextCov extracts architectural constraints from Agent Instructions that developers are already writing, lowering the barrier to architectural enforcement.
@@ 7. Limitations
We did not run controlled experiments with LLM agents because diffs generated by agents for specific repositories are not publicly available. While our detection results on existing codebases suggest ContextCov would prevent agent-introduced violations, a study comparing agent task success with and without ContextCov enforcement would directly demonstrate its value. Such an evaluation would require either access to proprietary agent execution logs or a large-scale study where agents are deployed on real repositories with ContextCov integration.
@@ 8. Conclusion
As agents take on more complex, multi-file tasks, certain classes of constraints become impractical to enforce through manual review. Style conventions, architectural boundary rules, and process workflow requirements are inherently difficult to verify by inspection. The specification that once resided in the developer’s mind must be externalized to Agent Instruction files, yet these files are passive text, and without automated enforcement, Context Drift accumulates silently.
ContextCov addresses this challenge by treating Agent Instructions as executable specifications, providing automated enforcement for constraints that are difficult to monitor manually. Through hierarchical constraint extraction, domain-specialized code synthesis, and multi-layer runtime enforcement, ContextCov transforms passive instructions into active invariants that prevent agent violations. Our evaluation on 723 open-source repositories demonstrates that ContextCov successfully extracts over 46,000 executable checks with 99.997% syntax validity and detects over 500,000 violations, with 81% of repositories containing at least one violation.
Future work could explore self-healing instructions where ContextCov suggests edits to Agent Instructions or code changes to restore compliance. Agent Instruction files could also adopt structured language features that enable deterministic check extraction without LLM involvement, improving reliability and reducing cost.
As autonomous agents become standard contributors to software repositories, frameworks like ContextCov will be essential for maintaining repository integrity, bridging the Reality Gap between what agents are told to do and what they actually do.

### 04 May 2026 v2

@@ 5. Discussion
The shift from human-supervised to autonomous agents fundamentally changes the role of project documentation. When agents produce changes that violate their instructions, this is not merely a technical bug but a symptom of a deeper mismatch between how we document intent and how agents consume it. We discuss the theoretical and practical implications of treating Agent Instructions as executable specifications.
@@ 5.2. The Agentic Feedback Loop
When an agent hits a ContextCov enforcement block, the error is fed back into its context window. This empowers the agent to reason about the failure. If the agent determines the rule was poorly scoped, it can autonomously propose a fix to the Agent Instructions. ContextCov transforms the agent from a passive consumer of documentation into an active maintainer of project specifications. Our design adopts a strict, fail-closed interpretation of documented constraints (Section 3.5). We argue that ambiguity in Agent Instructions is itself a defect. If a rule has valid exceptions, those exceptions should be specified explicitly. By blocking ambiguous cases, ContextCov forces Agent Instructions to become more precise over time.
@@ 5.3. Executable Interpretability
It is important to distinguish ContextCov from security sandboxes. Security mechanisms like SELinux (SELinux Project, [n. d.]), AppArmor (AppArmor, [n. d.]), and seccomp (Linux Kernel, [n. d.]) enforce mandatory access controls at the kernel level, while container runtimes like gVisor (Google, [n. d.]) and Firecracker (Amazon Web Services, [n. d.]) provide process isolation. Runtime verification frameworks (Havelund and Rosu, 2001; Chen and Roşu, 2007; Runtime Verification Inc., [n. d.]) monitor program execution against formal specifications. ContextCov is not designed to prevent a malicious actor from compromising a system. Rather, it is designed to prevent a well-intentioned but hallucinating agent (Lin et al., 2024) from degrading the repository. By restricting the execution environment through mechanisms like the Process Interceptor, ContextCov bounds the agent’s action space, ensuring that its exploratory problem-solving stays within approved architectural and operational limits.
@@ 6. Related Work
ContextCov draws on and extends several lines of prior research. Prior work has studied the “stale docs” problem (Aghajani et al., 2019), API documentation issues (Uddin and Robillard, 2015; Robillard and DeLine, 2011; Treude and Robillard, 2016), and README files (Prana et al., 2018; Gaughan et al., 2025). Techniques detect comment-code inconsistencies (Tan et al., 2007, 2012; Ratol and Robillard, 2017) and update comments after code changes (Panthaplackel et al., 2020; Liu et al., 2020), while requirements traceability research (Gotel and Finkelstein, 1994; Antoniol et al., 2002) maintains links between specifications and implementations. However, this prior work focuses on descriptive documentation (docstrings, comments) for human developers, whereas ContextCov targets prescriptive documentation (Agent Instructions) where stale docs become execution hazards since agents treat instructions literally.
@@
Research has also translated natural language to formal specifications (Ghosh et al., 2016; Brunello et al., 2019; Pandita et al., 2012; Zhong et al., 2009), and specification mining extracts properties from traces (Ammons et al., 2002; Gabel and Su, 2008; Kremenek et al., 2006; Ernst et al., 2007), while learning-based approaches infer linting rules (Bader et al., 2019). Tools like ESLint (esl, [n. d.]), Pylint (pyl, [n. d.]), Semgrep (sem, [n. d.]), Infer (Calcagno et al., 2015), and Error Prone (Aftandilian et al., 2012) enforce standards but require manual configuration or focus on general defects. ContextCov generates checks from Agent Instructions using domain-routed synthesis, leveraging code search techniques (Di Grazia and Pradel, 2023) to produce unified Python scripts operating on Tree-sitter ASTs, NetworkX graphs, and shell shims.
@@
Research on coding agents (Yang et al., 2024; Wang et al., 2025) has demonstrated autonomous repository navigation and issue resolution, with benchmarks like SWE-bench (Jimenez et al., 2024) and BigCodeBench (Zhuo et al., 2024) driving rapid progress. Agent alignment frameworks like Constitutional AI (Bai et al., 2022) and Guardrails (Guardrails AI, [n. d.]) focus on general-purpose safety constraints, while research on jailbreaking (Wei et al., 2024) reveals vulnerabilities in instruction-following. Recent work on detecting agent failures (Sharma et al., 2026; Barke et al., 2026) extracts behavioral rules from prompts and evaluates execution traces post-hoc for compliance violations. ContextCov complements such analysis by providing runtime enforcement: rather than detecting violations after execution, it intercepts non-compliant actions before they are finalized, enabling immediate self-correction.
@@
Finally, tools like Sonargraph (son, [n. d.]) detect dependency violations, reflexion models (Murphy et al., 2001) compare intended and actual architectures, Knodel and Popescu (Knodel and Daniel, 2007) survey architecture compliance checking approaches, and ArchJava (Chambers and Notkin, 2002) embeds architectural constraints directly in the programming language. These tools require architects to manually encode rules in tool-specific formats, whereas ContextCov extracts architectural constraints from Agent Instructions that developers are already writing, lowering the barrier to architectural enforcement.
@@ 7. Limitations
Several validity concerns apply to our evaluation. Generated Agent Instruction files may not perfectly match maintainer intent for repositories that lacked native instruction files, though we mitigate this by grounding generation in repository structure and conventions. Our false-positive analysis relies on a single LLM judge configuration (Claude Opus 4.5), which we partially address through direct human comparison on the same sample and by reporting agreement metrics. SWE-bench Lite includes only 12 Python-heavy repositories and 300 tasks, which may not reflect ecosystems with different languages, build systems, or instruction styles. While the 384-sample design supports 95/5 estimation for proportion metrics, some subgroup analyses may remain underpowered without stratified sampling, and our McNemar tests assume paired instances without correction for multiplicity across the three pairwise comparisons. Broader external validity would additionally require replication on other benchmarks and ecosystems beyond the 12-repository, 300-task Lite split.
@@ 8. Conclusion
As agents take on more complex, multi-file tasks, certain classes of constraints become impractical to enforce through manual review. Style conventions, architectural boundary rules, and process workflow requirements are inherently difficult to verify by inspection. The specification that once resided in the developer’s mind must be externalized to Agent Instruction files, yet these files are passive text, and without automated enforcement, constraint violations accumulate silently. ContextCov addresses this challenge by treating Agent Instructions as executable specifications, providing automated enforcement for constraints that are difficult to monitor manually. Through hierarchical constraint extraction, domain-specialized code synthesis, and multi-layer runtime enforcement, ContextCov transforms passive instructions into active invariants that prevent agent violations.
Our empirical evaluation on SWE-bench Lite demonstrates that relying on LLMs to self-police their compliance is both inefficient and error-prone. ContextCov outperforms both passive prompting and LLM reflection. As autonomous agents become standard contributors to software repositories, frameworks like ContextCov will be essential for maintaining repository integrity, bridging the gap between what agents are told to do and what they actually do.